# Halfway Escape Optimization: A Quantum-Inspired Solution for Complex Optimization Problems

### v2

@@ Abstract
This paper first proposes the Halfway Escape Optimization (HEO) algorithm, a novel quantum-inspired metaheuristic designed to address complex optimization problems characterized by rugged landscapes and high-dimensionality with an efficient convergence rate. The study presents a comprehensive comparative evaluation of HEO’s performance against established optimization algorithms, including Particle Swarm Optimization (PSO), Genetic Algorithm (GA), Artificial Fish Swarm Algorithm (AFSA), Grey Wolf Optimizer (GWO), and Quantum behaved Particle Swarm Optimization (QPSO). The primary analysis encompasses 14 benchmark functions with dimension 30, demonstrating HEO’s effectiveness and adaptability in navigating complex optimization landscapes and providing valuable insights into its performance. The simple test of HEO in Traveling Salesman Problem (TSP) also infers its feasibility in real-time applications.
@@
Optimization, Swarm intelligence, Metaheuristics, Halfway Escape Optimization(HEO).
journal: Applied Soft Computing
@@
1. we propose a new algorithm called HEO 2. we provide a comprehensive benchmark of 14 Objective Functions for our HEO algorithm 3. we evaluate the effectiveness of the new algorithm with the TSP
@@ 1 Introduction
The experimentation in this paper extends to the classical Traveling Salesman Problem (TSP), where the performance of the HEO algorithm is compared with that of Tabu Search and Random Search. TSP involves finding the shortest possible route for a salesman to visit a set of cities once and return to the starting city. Because TSP is a well-known NP-Hard Problem in the optimization realm, hence it is a well-testing problem with its exponentially increased time complexity[13]. The comprehensive analysis presented demonstrates the potential of the HEO algorithm as a powerful and adaptive optimization method. Overall, the paper first proposes HEO method and primary testing its effectiveness and convergence in 14 single objective functions.
@@ 2 Related Work
Particle Swarm Optimization (PSO) has gained widespread recognition as a bio-inspired metaheuristic algorithm rooted in the collective behavior of bird flocks. PSO was first proposed by Eberhart & Kennedy in 1995. PSO iteratively updates particle positions and velocities based on their individual and global best solutions, enabling effective exploration and exploitation of the search space[6]. While PSO has exhibited robustness in addressing various optimization problems, its potential for premature convergence and challenges in handling complex multimodal landscapes have been noted[24].
@@
The Grey Wolf Optimizer (GWO) draws inspiration from the hierarchical structure and hunting behavior of grey wolves in nature, which suggested by Mirjalili et al. in 2014. Through the alpha, beta, and delta wolf concept, GWO dynamically updates solution positions, showcasing promise in addressing complex optimization problems, particularly in multimodal landscapes. Nonetheless, the algorithm’s sensitivity to parameter settings and the need for domain-specific fine-tuning have been identified as areas for consideration Mirjalili et al..
@@ 4.1 Experiment Design
Table 1: Test Unimodal Functions
@@ 4.3 Experimental Comparison
Table 4: The Average Ranks of the Costs
For a more precise analysis, the paper conducts a straightforward rank aggregation procedure on the data presented in Table 3. Algorithms with lower associated costs are accorded higher ranks within this aggregation. The resulting aggregated rank data is displayed in Table 4. In Table 4, the Halfway Escape Optimization (HEO) achieves the highest rank across both unimodal functions (encompassing $F_{1}-F_{7}$) and multimodal functions (encompassing $F_{8}-F_{14}$). This suggests that HEO exhibits similar performance compared to the Grey Wolf Optimizer (GWO) specifically in handling multimodal Functions, it also demonstrates competitive performance relative to the other six algorithms tested in unimodal functions, even in scenarios involving relatively complex functions.
@@
5 Simple Test in Traveling Salesman Problem (TSP)
@@ 6 Discussion
The experimental results indicate that HEO outperforms other algorithms in terms of convergence speed and solution quality across a range of benchmark functions. The algorithm’s ability to balance exploration and exploitation, coupled with its adaptive strategies, has resulted in competitive performance in challenging optimization landscapes. The TSP test also validates its availability in path-planning applications.
@@ 7 Conclusion
Moreover, the adaptability of the HEO algorithm allows it to be applied to dynamic environments where the objective functions or constraints may change over time. By continuously adapting its search strategy and updating the particle positions, HEO can effectively track and respond to changing optimization landscapes, ensuring that high-quality solutions are maintained even in dynamic scenarios, including TSP applications.
Overall, the Halfway Escape Optimization algorithm has demonstrated its effectiveness and adaptability in addressing complex optimization challenges. Future work could focus on further refining the algorithm’s parameters and exploring its application in practical optimization problems. Additionally, the algorithm could be extended to address multi-objective optimization problems and dynamic environments or change the strategy of escape and skip, further enhancing its versatility and applicability.
@@
The experiment is data-available for replications after the relevant paper get published, provide the source code of HEO in: https://github.com/Spike8086/HEO.git, provide data of result and the notebook of experiment in: https://www.kaggle.com/code/spike8086/heo-formal-ver, and give the code for TSP test:https://www.kaggle.com/code/spike8086/heo-for-tsp
Symbols
Symbol Definition
$i$ The Current Number of Iteration
$N$ The Total Number of Iteration/Normal Distribution
$x_{i}$ The Position of The $i_{th}$ Quantum
$v_{l}$ The Global Velocity of The Quantum
$v_{g}$ The Local Velocity of The Quantum
$c_{i}$ The Escape Factor
$r_{1}$ The Random Factor for Escape Term
$r_{2}$ The Random Factor for The Whole Position Update
$r_{3}$ The Random Global Weight
$R$ The Radius for Scaling the Range of $r_{1}$
$f_{global}$ The Global Optimum of the Swarm
$f_{local}$ The Local Optimum of the Quantum
$A$ A Set Composed of the Positions of Two Quantum
$s$ The Position of Any Single Sample(Quantum/Particle)
$T$ Customized Function That Return the Optimum
$G$ Any Group
$S$ Any Set of Samples
$O$ Scaling Operation in HEO
$a_{i}$ Energy Level of Quantum
$r_{4}$ The Random Factor for Increment of Energy Level
$\sigma_{x_{i}}$ The Standard Deviation of Vector $x_{i}$
$b_{g}$ The Side Length/Radius of the Extra Constraint Area
$r_{5}$ The Random Factor for the Soft Edge
$b$ The Side Length/Radius of the Required Constraint Area
$\vec{r}$ The Random Factors for Random Skipping
$f^{*}$ The Function Value of the Optimum
$s^{*}$ The Solution of the Optimum
Abbreviation
Abbreviation Description
HEO Halfway Escape Optimization
TSP Traveling Salesman Problem
GA Genetic Algorithm
PSO Particle Swarm Optimization
ACO Ant Colony Optimization
VRP Vehicle Routing Problem
DE Differential Evolution
AFSA Artificial Fish Swarm Algorithm
FA Firefly Algorithm
QPSO/Q-PSO Quantum behaved Particle Swarm Optimization
GWO Grey Wolf Optimizer
SSA Salp Swarm Algorithm
SA Simulated Annealing

### v3

@@ Abstract
This paper first proposes the Halfway Escape Optimization (HEO) algorithm, a novel quantum-inspired metaheuristic designed to address complex optimization problems characterized by rugged landscapes and high-dimensionality with an efficient convergence rate. The study presents a comprehensive comparative evaluation of HEO’s performance against established optimization algorithms, including Particle Swarm Optimization (PSO), Genetic Algorithm (GA), Artificial Fish Swarm Algorithm (AFSA), Grey Wolf Optimizer (GWO), and Quantum behaved Particle Swarm Optimization (QPSO). The primary analysis encompasses 14 benchmark functions with dimension 30, demonstrating HEO’s effectiveness and adaptability in navigating complex optimization landscapes and providing valuable insights into its performance. The simple test of HEO in Traveling Salesman Problem (TSP), Pressure Vessel Design and Tubular Column Design infers its feasibility and potential weakness in real-time applications.
@@
Swarm intelligence, Optimization, Metaheuristics, Halfway Escape Optimization(HEO)
@@
we propose a new algorithm called HEO we provide a comprehensive benchmark of 14 Objective Functions for our HEO algorithm we further evaluate the effectiveness of the new algorithm with 3 classical engineering problems
@@ 1 Introduction
The experimentation in this paper extends to the classical Traveling Salesman Problem (TSP), where the performance of the HEO algorithm is compared with that of Tabu Search and Random Search. TSP involves finding the shortest possible route for a salesperson to visit a set of cities once and return to the starting city. Because TSP is a well-known NP-Hard Problem in the optimization realm, hence it is a well-testing problem with its exponentially increased time complexity[15]. The comprehensive analysis presented demonstrates the potential of the HEO algorithm as a powerful and adaptive optimization method. Overall, the paper first proposes HEO method and primary testing its effectiveness and convergence in 14 single objective functions.
@@ 2 Related Work
Particle Swarm Optimization (PSO) has gained widespread recognition as a bio-inspired metaheuristic algorithm rooted in the collective behavior of bird flocks. PSO was first proposed by Eberhart and Kennedy in 1995. PSO iteratively updates particle positions and velocities based on their individual and global best solutions, enabling effective exploration and exploitation of the search space[7]. While PSO has exhibited robustness in addressing various optimization problems, its potential for premature convergence and challenges in handling complex multimodal landscapes have been noted[27].
@@
The Grey Wolf Optimizer (GWO) draws inspiration from the hierarchical structure and hunting behavior of grey wolves in nature, which suggested by Mirjalili et al in 2014. Through the alpha, beta, and delta wolf concept, GWO dynamically updates solution positions, showcasing promise in addressing complex optimization problems, particularly in multimodal landscapes. Nonetheless, the algorithm’s sensitivity to parameter settings and the need for domain-specific fine-tuning have been identified as areas for consideration Mirjalili et al.
@@ 4.1 Experiment Design
[…]
Table 2: Test Multimodal Functions
@@ 4.3 Experimental Comparison
[…]
For a more precise analysis, the article performs a straightforward rank aggregation procedure on the data presented in Table 3. Algorithms with lower associated costs are accorded higher ranks within this aggregation. The resulting aggregated rank data is displayed in Table 4. In Table 4, Halfway Escape Optimization (HEO) achieves the highest rank in both unimodal functions (encompassing $F_{1}-F_{7}$) and multimodal functions (encompassing $F_{8}-F_{14}$). This suggests that HEO exhibits similar performance compared to the Grey Wolf Optimizer (GWO) specifically in handling multimodal Functions, it also demonstrates competitive performance relative to the other six algorithms tested in unimodal functions, even in scenarios involving relatively complex functions.
Table 5: Search Time(s/1000 iters) of the Algorithms
@@
5 Traveling Salesman Problem (TSP)
@@
6 The Pressure Vessel Design Problem
The pressure vessel design is a classical engineering problem, the objective is to change the structure of the vessel to reduce its manufacturing cost [22]. The optimization algorithm tested in this problem could well validate its performance in complex constrained situations. As Fig.14 and Fig.15 shown, the vessel composed by a cylinder shape of the structure welded with a sphere bottom. By changing from $x_{1}$ to $x_{4}$ values($x_{1}$:($R$-Radius of the Shell), $x_{2}$:($L$-Length of the Shell), $x_{3}$:($T_{s}$-Thickness of the Shell) ,$x_{4}$:($T$-Thickness of the dish end)), the task is to minimize the cost in equation(26).
Figure 14: The Structure of the Vessel [14]
Figure 15: The Heatmaps of the Vessel [18]
$Consider:$
$\vec{x}=[x_{1},x_{2},x_{3},x_{4}]$
$Minimize:$
$f(\vec{x})=0.6224x_{1}x_{2}x_{3}+1.7781x_{1}^{2}x_{3}+3.1161x_{2}x_{4}^{2}+19.84x_{4}x_{1}^{2}$
$Subject\quad to:$
$g_{1}(\vec{x})=-x_{1}+0.0193x_{3}\leq 0$
$g_{2}(\vec{x})=-x_{3}+0.000954x_{3}\leq 0$
$g_{3}(\vec{x})=-\pi x_{3}^{2}x_{4}+\frac{4}{3}\pi x_{3}^{3}+1296000\leq 0$
$g_{4}(\vec{x})=x_{4}-240\leq 0$
$Variable\quad Range:$
$0.00625\leq x_{1}\leq 1.25$
$0.00625\leq x_{2}\leq 1.25$
$10\leq x_{3}\leq 200$
$10\leq x_{4}\leq 200$
Table 7: Pressure Vessel Design in 30 times Validation
PSO GWO HEO GA QPSO
$\mu_{cost}$ 9938.35 9197.10 8638.74 8757.88 7680.54
$\sigma_{cost}$ 1952.30 1890.72 1268.32 1533.92 825.02
$T_{s}^{*}$ 0.8125 1.25 1.25 1.0 0.9375
$T_{h}^{*}$ 0.4375 0.5625 0.5625 0.5 0.5
$R^{*}$ 40.3227 58.9299 57.6567 51.4577 88.9923
$L^{*}$ 200.0 40.4031 47.4311 48.3081 112.6784
[…]
As Table 7 and Fig.16 illustrated, the HEO algorithm show a relatively well small cost among the others, but larger than the cost of QPSO. The Pressure Vessel Design Problem is a linear combination problem with constraints, which is similar to the case of the Step Function. From the HEO’s cost in the Step Function(or $F_{2}$) being larger than QPSO,it is not unexpected that HEO perform worse than QPSO in this problem, which might indicate HEO’s weakness when solving the complex problem with simple constraints.
Figure 16: The Distribution of Cost for Pressute Vessel Design
7 Tubular Column Design Problem
As Fig.17 shown, the Tubular Column this problem is structural part that is a cylindrical shaft, which may be solid, or made of a single circular cross-section and produced from metal, concrete and other materials [1]. It is commonly used within constructing projects where it is required to support beams, as well as other elements, such as columns, and so on, and it is also used in bridges and other types of structures.
Typically, tubular columns are superior and efficient compared to solid columns because they can resist torsion, bending, and shear force. The ideal is to minimize the cost of realising the column which is captured by the variables d, the average diameter of the column in centimeters and t the thickness of the column in centimeters.
In Table 8, the paper set 5 algorithms with same parameter as the previous tests in the experiment(HEO with $c_{m}ax=1$), the AFSA is do not show any inclination of converge for testing engineering problems, hence it get rid from it. The HEO has the smallest mean value of cost compared to other algorithms in 30 times validations, might illustrate its good performance in the tubular column problem,or relatively simple function with complex constraints. As for the deviations, the GA get the smallest one,indicating its stability to some extent.
Figure 17: The Structure of the Tubular Column Design[1]
$Consider:$
$\vec{x}=[x_{1},x_{2}]$
$Minimize:$
$f(\vec{x})=9.8x_{1}x_{2}+2x_{1}$
$Subject\quad to:$
$g_{1}(\vec{x})=\frac{P}{\pi x_{1}x_{2}\chi_{y}}-1\leq 0$
$g_{2}(\vec{x})=\frac{8PL^{2}}{\pi^{3}Ex_{1}x_{2}(x_{1}^{2}+x_{2}^{2})}-1\leq 0$
$g_{3}(\vec{x})=\frac{2}{x_{1}}-1\leq 0$
$g_{4}(\vec{x})=\frac{x_{1}}{14}-1\leq 0$
$g_{5}(\vec{x})=\frac{0.2}{x_{2}}-1\leq 0$
$g_{6}(\vec{x})=\frac{x_{2}}{8}-1\leq 0$
$Variable\quad Range:$
$2\leq x_{1}\leq 14$
$0.2\leq x_{2}\leq 0.8$
Among them in equation(27), $P=2300$ means compressive load with unit $kg_{f}$, $chi_{y}=450$(is yield stress, has unit $kg_{f}/cm^{2}$), $E=0.65*10^{6}$ is elasticity,unit is $kg_{f}/cm^{2}$),as for $L=300$, is length of column with unit $cm$.
Table 8: Tubular Column Design in 30 times Validation
PSO GWO HEO GA QPSO
$\mu_{cost}$ 32.105096 32.147502 31.704160 32.224944 32.144308
$\sigma_{cost}$ 0.419599 0.366717 0.760805 0.015721 0.369621
$d^{*}$ 7.1031 7.1013 7.1477 8.1346 7.1013
$t^{*}$ 0.229 0.2295 0.2279 0.2 0.2292
[…]
@@ 8 Discussion
The experimental results indicate that HEO outperforms other algorithms in terms of convergence speed and solution quality across a range of benchmark functions. The algorithm’s ability to balance exploration and exploitation, coupled with its adaptive strategies, has resulted in competitive performance in challenging optimization landscapes. The TSP and test in Tubular Column Design also validates its availability in path-planning applications.
@@ 9 Conclusion
Moreover, the adaptability of the HEO algorithm allows it to be applied to dynamic environments where the objective functions or constraints may change over time. By continuously adapting its search strategy and updating the particle positions, HEO can effectively track and respond to changing optimization landscapes, ensuring that high-quality solutions are maintained even in dynamic scenarios, including various engineering applications.
Overall, the HEO has demonstrated its effectiveness and adaptability in addressing complex optimization challenges. Future work could focus on further refining the algorithm’s parameters and exploring its application in practical optimization problems. Additionally, the algorithm could be extended to address multi-objective optimization problems and dynamic environments or change the strategy of escape and skip, further enhancing its versatility and applicability.
@@
The experiment is data-available for replications, provide the source code of HEO in: https://github.com/Spike8086/HEO.git, provide data of result and the notebook of experiment in: https://www.kaggle.com/code/spike8086/heo-formal-ver, and give the code for TSP test:https://www.kaggle.com/code/spike8086/heo-for-tsp